I have this problem to do in my math end of term exam revision. I do not understand how to work it.
The problem is: in what year is the baseball card worth 20$? and the equation is: V(T)= 4.75(1.12)^t ...t=0 is year 1980.
Do I try different years for the fun of it or is there a way to calculate it?
Thanks for the help!

Answer by stanbon(48535)

The problem is: in what year is the baseball card worth 20$? and the equation is: V(T)= 4.75(1.12)^t ...t=0 is year 1980.
---------------------
Let V(T) be $20 and solve for "t":
20 = 4.75(1.12)t
1.12^t = 4.2105
Take the log of both sides to get:
t*log(1.12) = log 4.2105
t = [log4.2105]/[log(1.12)]
t = 12.6851 years
-------------------------
If t=0 is 1980 you add 12.6851 years and you are
in the year 1993.
Cheers,
Stan H.